OpenStudy (anonymous):

Rewrite y''+y'+y=0 as a system of first order differential equations. Find the eigenvalues and eigenvectors of the system.

OpenStudy (anonymous):

\[\lambda^{2} + \lambda + 1 =0\] solve it

OpenStudy (anonymous):

Is it \[\lambda= +/- \sqrt{-3/4}-1/2 \] ?

OpenStudy (anonymous):

correct those are the eigenvalues \[-\frac{ 1 }{ 2 }\pm \frac{ i \sqrt{3} }{ 2 }\] can you put them in sin and cos form to get the eigenvectors

OpenStudy (anonymous):

I'm not sure, could you walk me through it?

OpenStudy (anonymous):

are the sin and cos forms \[e ^{-1/2t}(\cos \sqrt{3}/2t + i \sin \sqrt{3/2}t)\] and \[e ^{-1/2t}(\cos \sqrt{3}/2t - i \sin \sqrt{3/2}t)\] ?

OpenStudy (anonymous):

\[y(t) = e^{-1/2t}(c1*\cos(\sqrt{3}/2)+c2*\sin(\sqrt{3}/2))\]

OpenStudy (anonymous):

ok, i think I see that

OpenStudy (anonymous):

Is y(t) our eigenvector here?

OpenStudy (anonymous):

yup site is too slow!

OpenStudy (anonymous):

lol, I had to wait an hour earlier this morning to get on. Thanks for your help, I really appreciate it :)

OpenStudy (anonymous):

no problem. btw are you engineer?

OpenStudy (anonymous):

yes, mechanical. I'm a 3rd year.

OpenStudy (anonymous):

im electrical :)

OpenStudy (anonymous):

cool :) what year? or have you graduated already?

OpenStudy (anonymous):

my 2nd yr in masters